PVFF Action Plan
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
View only
 
 
ABCEFGHIJKLM
1
ID
Relates to ID
Date Raised
Parish
Assigned To (Organisation)
Assigned To (Person)
Original Due Date
Latest Due Date
Action
Last Updated
StatusStatus Notes
2
3201/01/2014
Stanford Dingley
TWTW Stanford Dingley strategy
Determine location of ingress to sewer network in Stanford Dingley.
22/04/2018In progressDepth logger data has been reviewed but  not conclusive for points of ingresss. TW have confirmed they will now be trying smoke testing.

24/01/2017 work to line a sewer to be completed by 02/2017. However there remain problems in SD that require regular tanker visits and thus TW are carrying out further investigations via manhole lifting. 2 of 4 loggers are still working.

18/5/2017 Meeting held in SD with BC, MH, KL and Andrew Hagger TW, Denise Kinsella TW and James Hoar TW.Maps of sewage pipework discussed and it was agreed that two flow monitors near Frilsham that were key to locating point of ingress were not working properly and possibly needed replacing. One data logger on Jennetts Hill to be moved up to Frilsham area to triangulate location of ingress. Some manholes still needing to be replaced on Jennetts Hill. An impermeable area survey was due to be done in Frilsham in approx 2-3 weeks time. It was noted that it may still take some time to locate point of ingress and it may not be possible to do so until next winter and return of wetter weather. TW were keen to get to bottom of problem as SD was one of only two areas that had confounded them.

23/05/2017 TW to carry out an IAS survey when we have a wet winter. New Flygt pumps to be implement this summer in SD station.

26/09/2017 Flygt pumps installed July 2017 (and a 2nd in Nov 2017) but damage done to 'ducks foot' base during installation. Repair due to start 26th Oct and due to take two days. All manhole covers on Jennetts Hill have been replaced and the impermeable survey has been completed showing no misconnections in the Frilsham area, which further narrows the area of search. The data logger is still to be moved around but there is a wider programme of depth monitoring in the Thames Valley area so we are looking to move these but also see if we can install some flow monitors/data loggers that feed directly into our modelling software. We have also completed flow monitoring on the rising main in June which gave us information on the pump rate through the rising main.

23/01/2018 Ducks foot (bases for the pumps) have been replaced and thus second Flygy pump installed. Still trying to narrow where the ingress is happening. Data loggers are still to be moved. Track to sewer pumping station needs attention

22/04/2018 The data loggers were moved at the end of February and the information from the recent heavy rainfall events in March/early April is being looked at to see what any next steps may be. Map with location of manholes where loggers employed sent to SD PC/FW.
3
454115/09/2015
All Parishes
PVFF01/10/2015Devise a village-centric account of causes, consequences and escalating risk levels. Mapping stage 212/01/2016On holdThis activity should be phase 2 of the Pang Valley flood risk mapping exercise
4
4815/09/2015West IlsleyWBCS Clark01/03/201631/03/2016West Ilsley to East Ilsley Road - investigate new road crossing at EA pump station.13/04/2018In progressTo be investigated in the 15/16 FY. No progress yet as of 12/01/2016.

20/09/2016 update: A pipe has been found. WBC to investigate opening up the pipe, liaising with the landowner on either side. This could not be done in FY16/17 as budget is committed.

24/01/2017 SC aims to look for funds in 17/18. PVFF Tech group will include this area in their slow sections.

23/05/2017 Pending suitable funds and staff availability.

26/09/2017 Currently on hold

13/4/2018 WBC Aim to complete this year (FY 18/19) subject to staffing (Graduates are being interviewed).
5
5015/09/2015West IlsleyWBCS Clark01/03/201631/03/2016WBC West Ilsley drainage scheme.
Capital investigation into the upgrade of the surface water system through West Ilsley (Defra funded feasibility)
13/04/2018In progressSite meeting held with consultant on 24th November. Feasibility has been completed. Level survey to be completed. To progress further SC must meet consultant on the cost of feasibility.

24/01/2017 This study is in progress and study should complete by end of 03/2017.

23/05/2017 Feasibility study in progress.

26/09/2017 Business case submitted to the EA. It is likely that the recommendation will be for property level protection.
13/04/2018 The business case was submitted to the EA, which concluded that the scheme is not financially viable. The only option is for Property Level Protection (PLP), for which funding is available. FW to establish how many properties are eligible for PLP, but do not have it.
6
51.15112/01/2016West IlsleyTW01/01/2017TW West Ilsley strategy.
TW to create a whole drainage strategy for West Ilsley
22/04/2018In progressWest Ilsley Drainage strategy is within East Ilsley Drainage Strategy as the name is based on the location of the sewage treatment works. Stage 1 is on the TW website. Stage 2 due 1/9/2016

23/05/2017 Survey conducted in Compton but then it was dry. Not the best data. Awaiting wetter weather.

26/09/2017 Groundwater remains low. No progress.

23/01/2018 With regards to the update / holding reports, these have been delayed, the aim is now to have them published in February 2018 Apologies for the delay.

22/04/2018 The reports are still with TW communications team finalising the last bits around tone etc. They are hopeful these will be available before the meeting on 26th April, or very soon after.
7
5815/09/2015
Hampstead Norreys
WBCS Clark01/03/2016WBC Hampstead Norreys scheme.
Investigate a permanent overflow channel around Hampstead Norreys as provided during the flooding.
13/04/2018In progress1. Overflow swale in field
2. Blockage on bridge on B4009(underground channel to water meadow) – this wasn’t part of the WBC bid and there is a restriction below the B4009
3. Sewage ingress was missed. WBC resubmitted a project mandate for the works - pending a response.

23/05/2017 Preliminary Design and Business Case 2017/18, Detailed design 2018/19, construction 2019/20

26/9/2017 Preminary design and Business Case in progress

23/01/2018 Still in progress.
08/04/2018 Business Case submitted. EA now checking. Grant allocation for detailed design submitted this year. EA have given funds to the detailed design phase which will be done this year.
8
5915/09/2015
Stanford Dingley
WBCS Clark01/03/2016WBC Stanford Dingley bypass scheme.
Bid submitted to Defra (RFCC Grant Funding) for permanent bypass channel around Stanford Dingley.
08/04/2018In progressBid submitted through EA, awaiting confirmation of RFCC programme. Funding awarded for the investigation and feasibility in the 2015/16 financial year. UPDATE - The estimated cost of the study is higher than the grant allocation. Application for additional grant money submitted to the EA. It is Surface Water flooding. WBC have found money to conduct the study, design and create the PAR (aka the business case). PAR work has started with consultant working on this.

20/09/2016: when the consultants have a design and a model, SC to share with Stanford Dingley (Mark) by end of 03/2017

23/05/2017 PAR would allow 75 year protection. Business case stacks up.

26/09/2017 Detailed design in progress.

23/01/2018 WBC agents (Dalcour Maclaren) have visited the village. Next steps should involve a topographical survey. Indicated that diggers might arrive this year (2018).
08/04/2018 Current cost/benefit ratio in Business Case is not sufficient to take the scheme forward. SC is trying to tweek the design in order to improve the ratio. A lot depends on the cost of diverting utilities under the road so he plans to do a trial hole next month to determine what needs doing and how much it will cost.
9
6612/01/2016ComptonTW01/01/2017TW Compton strategy.
TW to create a whole drainage strategy for Compton
22/04/2018In progressCompton Drainage Strategy Stage 1 on TW website

20/09/2016: TW will implement flow monitors in Compton to see if there is infiltration. Also planning to do CCTV work. Depth monitors in place, also in Hampstead Norreys.

24/01/2017 Update is the same for 20/09/2016 - current low water levels prevent data gathering. TW to share trigger level for groundwater level when leaks might start into the sewer system. Installing rising main monitors East Ilsley pumping station and HN pumping station (budget is committed).

23/05/2017 CCTV surveys scheduled, but need more groundwater for the work to succeed.

26/09/2017 CCTV surveys scheduled, but need more groundwater for the work to succeed. TW will create a holding document update for WI, EI, Co, HN assessments.

23/01/2018 3 Flow Monitors were installed in Compton from the 24/03/2017 to the 08/05/2017 (that have now been removed), during a fairly dry period, as such no signs of infiltration were seen / expected. This is now available as a good baseline for when the higher ground water levels return. With regards to the update / holding reports, these have been delayed, the aim is now to have them published in February 2018. TW to send link when reports are published to PVFF.

22/04/2018 The reports are still with TW communications team finalising the last bits around tone etc They are hopeful these will be available before the next meeting on the 26th April , or very soon after.
10
78.27823/01/2017
Pangbourne
EADMcGPangbourne and Tidmarsh flood modelling.
EA to share results of the CH2M modelling for the Pang and Sulham Brook in the Pangbourne and Tidmarsh areas.
22/04/2018In progressMeeting to be arranged with Pangbourne project team re modelling CH2M

26/09/2017 Meeting held on 7th July in Pangbourne with site visit including Moor Copse and Tidmarsh. Notes from meeting are on website. Further photos and information subsequesntly sent to CH2M to assist with their modelling. Update awaited. Phase 1 modelling by Ch2m has been updated with new information provided by PVFF people - this now matches what happened in 2007 and 2014. This means that EA will fund further modelling.
Option then to be considered:
1. Home Farm bypass
2. Flood Storage area
3. Raise banks along the Pang

Review meeting to be arranged after stage 2 modelling is complete.

23/01/2018 Rashmi Krishnan is new Project manager for this. CH2M have successfully managed to create a flood model that represents what happened in 2007. This took longer than expected. A non-technical reason for this delay is because they couldn’t adjust one of the gauging stations to reflect what happened in 2007. I think this was the Tidmarsh gauge. This has now been resolved. The next step is to update the hydrology to show the changes that have happened since 2007. These are the things such as the flood wall, channel maintenance and the other improvements that were shared with us at the site meeting back in the Summer. The consultants are going to inform us how they plan to do this in January. We can then identify if anything has been missed.

23/01/2018 EA to invite PVFF Pangbourne and Tidmarsh reps to review the model and options for a scheme.

22/4/2018 Since the last meeting, the Environment Agency received the outputs from the updated modelling. This modelling is now correctly representing the 2007 and 2014 events.
• The modelling shows that there are now much fewer properties at risk because of the works done between 2007 and 2014. Therefore, out of the flood alleviation options considered, there were no economically viable options to tackle the remaining risk.
• However ,there are still a significant number of residual properties at risk. Therefore, the Environment Agency have asked CH2M to consider if any smaller scale options could reduce the risk, even if they could not eradicate the risk. We are still waiting for a response on this.
• KL requested a meeting to review the research to date, understand which properties are still at risk, see what is still being considered. DM has spoken to Rashmi and she is going to contact CH2M to ask them for their availability to meet with representatives of the PVFF. She expects to hear back from them very soon and schedule in the meeting in the near future.
11
7924/01/2017
Pangbourne
WBCPangbourne mill clearing.22/04/2018On holdThe build up of slit downstream of the mill has been removed. A budget estimate to clear and remove silt immediately upstream of the mill would be about £5k, plus another £4k to replace the trash screen.

23/05/2017 It may be that further work is not needed. Initial clearing may have helped. Pangbourne PC to observe.

26/09/2017 WBC to show Thames Water (mill owners) where extra silt removal should be done upstream.

23/01/2018 No progress.


22/04/2018 SB and KL are happy that in the recnt Flood Alert over Easter the river was able to manage with the amount of silt in situ. The current operation of the mill sluices is allowing more water down the mill side and this will be encouraging the silt to be drawn downstream anyway. Current decision is to leave until further works planned on TW site, possible to look at reducing obstruction to fish passge, when this may be re-evaluated.
12
8024/01/2017
All Parishes
WBCJacobs report - impact of climate change.
WBC to investigate if the Jacobs report might change if the rain levels were increased by 30%
18/05/2017On hold18/05/2017 No progress yet.
23/05/2017 Put on hold until National Goundwater study is published.
23/01/2018. Peter to check if this has been published.
13
8224/01/2017
All Parishes
PVFFG DavidsonPVFF Flood Warden leaflet.
Draft a flood warden role document to circulate across the forum
22/04/2018In progress26/09/2017 GD has created a draft. Now all PVFF reps should review.

21/03/2018 WBC have been included and the updated document is now published on the PVFF website (and linked to from the homepage). Kay and Stephen to tweak and make extra user friendly.

22/04/2018 No further work done at present
14
8424/01/2017TidmarshPVFFTidmarsh bank strengthening.
Bank strengthening work was done. The work was not done well and needs to be completed and to be fenced off. Kay to raise with them.
26/04/2018Pending SignOffCL has checked the re-instated bank 20/5/2017 and happy that it is now re-vegetating and it also appears to be partly fenced off by BBOWT staff. However at the lasher weir the trash screen is still missing.  CL to follow up with BBOWT.  The logs on the bank also need to be secured with hoops of rebar.  GD and CL will do this.

23/01/2018 GD and CL still to complete.

26/04/2018 Both have booked to do this over weekend of 12/13th May
15
87.18723/01/2017
All Parishes
PVFFPeter McGeehinGroundwater project.
PM to seek written confirmation on whether WBC, TW will commit funding to assist further research beyond the MSC (as EA (DB) confirmed that they would if the others did).
22/04/2018In progressMSC is due to finish on 23/08/2017. PM to receive the MSC thesis that can then be shared.

21/01/2018 MSc Thesis complete. Copy sent to PMcG and Tech Sub Group.

22/04/2018 Meeting held early 2018 with EA, TW and PVFF. Support for taking project further. Decision made that a further 2 MScs are justified. PM has approached a number of Universities with two project briefs. Much interest shown However, we are too late for some institutuions as their students will have already chosen their projects. We have heard that Birmigham have a student who will take the project. We are well placed with the other Universities for our second project to be selected later in the year.
16
8926/09/2017
Pangbourne Parish Council
John HiggsThames Path.
Kay to contact WBC (Carolyn) and follow up whether the Thames Path work has been completed.
23/01/2018In progress23/01/2018 Some work done. JH to and PPC to liaise with National Trails Officer and Elaine Cox at WBC if further work needed.
17
9326/09/2017PVFFAll parishesParish Donations.
Each PVFF rep to ask their local parish council to donate a recurring amount of £100 per annum
22/04/2018In progress23/01/2018 £100 received from Compton PC. Emails have been sent to each parish council.

22/04/2018 £100 donations received from East Ilsley, Compton, Bradfield and Tidmarsh. Pledged donations from Pangbourne (arriving May) and Bucklebury.
18
9426/09/2017PVFFAll parishes2007 and 2014 reports.
Each PVFF rep to look at the 2007 and 2014 reports for their parish to see what was recommended.
22/04/2018In progress23/01/2018 KL provided spreadsheets for each parish, with their recommendations and a request for comments and notes as to whether completed. Stuart and Carolyn will now review and revert to the Flood Forum with an update as to why recommendations not completed, whether they have been superceded or whether they can be added to the Action Plan.

22/04/2018 New member of team (Helen) at WBC updating notes on recommendations. Hope to receive prior to meeting on 26/4/2018
19
9623/01/2018EAD. DukesSluice management in emergencies.
Does the EA have contact details for sluice / water management structures on the Pang which might be needed in emergency? EA to confirm with its emergency control room what information it needs and how it would access it.
22/04/2018In Progress22/4/2018 The Environment Agency do have contact details for key features on the Pang, which they have procedures to contact in emergencies.
These are:
o Tidmarsh Mill and Sluice – Requested to open sluice gate on Flood Alert
o Purley Hall Penstock – Requested to open outlet sluice on Flood Alert when lake reaches 300mm freeboard
o Pangbourne Mill and Sluice and Thames Water Site Pangbourne – Thames Water are automatically alerted when Flood Alert is reached. EA have a specific contact in case of issues.
o Bucklebury FAS – Specific contact in case of issues
o Compton Debris Screen – Specific contact in case of issues
o Briars Close area, Pangbourne – Specific contact in case of issues
20
9723/01/2018PVFFP AllisonUpdate the sluice operating procedures are in a consistent form, based on triggers23/01/2018New
21
9923/01/2018EARiver maintenance results.
Each year (in Jan) EA to present what maintenance was actually done in the previous year (with a MAP)
22/04/2018In progress22/4/2018 Apologies for the delay, caused by EA staff changes. This should be provided soon, hopefully prior to the meeting on 26th April.
22
10223/01/2018ComptonEAEA to investigate possible further watercourse improvements downstream of Aldworth Road22/04/2018In progress22/04/2018 Environment Agency has been talking with Dr Peter McGeehin, however, unfortunately, the issues were not able to be solved this way. Therefore, the Environment Agency has contacted the riparian owners directly.
23
10423/01/2018ComptonPVFFPeter McGeehinCompton Channel Maintenance.
Peter to contact to Yattendon Estate re riparian responsibilities.
23/01/2018New22/04/2018 The Environment Agency has contacted the riparian owners directly.
24
10523/01/2018
Hampstead Norreys
EAEA to inspect downstream of the B4009 bridge to ensure that the river channel is not being restricted by garden extensions.22/04/2018In progress22/04/2018 Environment Agency field team have been requested to go to site. The results of their inspection should be available before the meeting on 26th April.
25
26
27
28
29
30
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
60
61
62
63
64
65
66
67
68
69
70
71
72
73
74
75
76
77
78
79
80
81
82
83
84
85
86
87
88
89
90
91
92
93
94
95
96
97
98
99
100
Loading...
 
 
 
All Open Actions
All Actions (this page is the one to edit)
WBC Open Actions
TW Open Actions
EA Open Actions
PVFF Open Actions
Update Log
Data
 
 
Main menu